Determine the derivative of "y = e"arcsin"^x"


1
Expert's answer
2021-06-12T05:16:13-0400

"y=e^{sin^{-1}x}"

Differentiating with respect to x

"\dfrac{dy}{dx}=e^{sin^{-1}x}\times\dfrac{1}{\sqrt{1-x^2}}\times1"

"y'=\dfrac{e^{sin^{-1}x}}{\sqrt{1-x^2}}"